# Problem-Solving Logic Guided Curriculum In-Context Learning for LLMs Complex Reasoning

### v1

@@ Abstract
In-context learning (ICL) can significantly enhance the complex reasoning capabilities of large language models (LLMs), with the key lying in the selection and ordering of demonstration examples. Previous methods typically relied on simple features to measure the relevance between examples. We argue that these features are not sufficient to reflect the intrinsic connections between examples. In this study, we propose a curriculum ICL strategy guided by problem-solving logic. We select demonstration examples by analyzing the problem-solving logic and order them based on curriculum learning. Specifically, we constructed a problem-solving logic instruction set based on the BREAK dataset and fine-tuned a language model to analyze the problem-solving logic of examples. Subsequently, we selected appropriate demonstration examples based on problem-solving logic and assessed their difficulty according to the number of problem-solving steps. In accordance with the principles of curriculum learning, we ordered the examples from easy to hard to serve as contextual prompts. Experimental results on multiple benchmarks indicate that our method outperforms previous ICL approaches in terms of performance and efficiency, effectively enhancing the complex reasoning capabilities of LLMs. Our project will be publicly available subsequently.
@@ 1 Introduction
Large language models (LLMs) Ouyang et al. (2022); Ye et al. (2023); Bahrini et al. (2023) can rapidly acquire new capabilities through in-context learning (ICL) to solve many new tasks Wies et al. (2024); Xu et al. (2024), and can be extended through chain of thought (CoT) Wei et al. (2022) to solve many tasks that require complex reasoning Hao et al. (2023); Zhang et al. (2023). Researchers believe that through ICL, LLMs can implicitly learn the problem-solving patterns demonstrated in contextual examples and apply them to new tasks Bhattamishra et al. (2023); Dai et al. (2023). This means that LLMs have the ability to learn and apply problem-solving patterns on the spot from given examples.
In recent years, supervised fine-tuning (SFT) methods Dong et al. (2023) and reinforcement learning optimization reasoning methods Du et al. (2023); Guo et al. (2025) have been able to significantly enhance the reasoning abilities of LLMs through training. Despite this, due to the unique characteristic of ICL that it can enhance problem-solving capabilities without training, it still holds value as significant as the methods mentioned above, especially when facing the need to reduce costs or quickly apply to new tasks. Relevant work Hsieh et al. (2023) has already shown that LLMs possess a wealth of basic knowledge and fundamental capabilities that can be effectively activated through a small number of examples. Particularly, LIMO Ye et al. (2025) fine-tuned a large language model with only a few hundred examples and achieved results that are close to or even on par with the current state-of-the-art reinforcement learning optimization inference. Therefore, we believe that the ICL capabilities of current LLMs are still far from being fully realized. There is a need to design better prompts to effectively enhance the effectiveness of ICL.
@@
In this work, we innovatively propose an problem-solving logic guided curriculum ICL method, which constructs the optimal ICL prompt for the query based on problem-solving logic. The Question Decomposition Meaning Representation (QDMR) Wolfson et al. (2020) decomposes complex problems into several sub-questions for solving and formalizes these sub-questions with 13 custom "operations", which we refer to as problem-solving logic. Figure 1-(a) shows an example of problem decomposition and transformation into problem-solving logic. Although it cannot directly solve the problem, the problem-solving logic describes the steps required for solving and the order of these steps in formal language. Therefore, it can accurately measure the intrinsic connections between examples and construct a sequence of demonstration examples that are conducive to problem-solving. Figure 1-(b) shows an example of curriculum ICL. We select examples with similar problem-solving logic, which can help LLMs learn how to solve similar problems. Subsequently, we measure the difficulty of these examples by the number of problem-solving steps. The greater the number of steps, the more reasoning steps are involved, meaning the problem is more difficult to solve. Relying on the principles of curriculum learning, we order these examples from easy to hard to serve as the final in-context prompt.
@@ 3.2.2 Demonstration Examples Ordering
Random Random Random 76.5% 46.5% 73.8% 75.8% 65.1% 67.53%
VoteK KNN Similarity 74.9% 44.9% 76.7% 75.4% 69.0% 68.19%
PromptSO PCA Eigenvalue 77.3% 43.7% 77.7% 75.6% 67.7% 68.40%
AutoCoT K-means Similarity 77.5% 47.2% 75.3% 76.0% 71.2% 69.44%
CoT + Fewshot Fixed Fixed 80.5% 44.5% 79.4% 75.1% 69.4% 69.79%
SA-ICL KNN Entropy 78.8% 47.6% 77.9% 78.5% 66.8% 69.95%
AL-ICL KNN Similarity 80.8% 45.7% 78.2% 77.9% 68.1% 70.13%
Ours PSL Curriculum 83.4% 50.8% 81.1% 75.0 71.6% 72.37%
Table 1: The table presents a comparison of experimental results across different benchmarks using Llama3-8B, demonstrating the accuracy contrast between various ICL methods. Avg represents the average accuracy across the different benchmarks. The best and second-best performances are highlighted in bold and underlined, respectively.
@@ 4.1 Experimental Setup
We evaluate the effectiveness of our method on the Llama3-8B model. For each benchmark, we select demonstration examples from its training set to form prompt information to evaluate each test set data. For the SVAMP dataset, we adopted the same evaluation strategy as in previous work Patel et al. (2021), using ASDiv-a Miao et al. (2020) and MAWPS Koncel-Kedziorski et al. (2016) together as the training set. To ensure a fair comparison, the number of selected examples is based on the settings in CoT Wei et al. (2022) for different benchmarks, and our experiments do not exceed that limit.
@@
Prioritize simplicity w/ order 82.3% 47.6% 79.5% 75.5% 69.0% 70.79%
w/o order 82.5% 47.2% 78.8% 76.1% 68.1% 70.55%
Prioritize difficulty w/ order 81.8% 44.9% 77.9% 76.6% 67.7% 69.77%
w/o order 81.6% 46.1% 79.6% 77.0% 67.2% 70.29%
Select Randomly w/ order 81.3% 50.6% 80.2% 76.1% 70.3% 71.70%
w/o order 80.9% 48.6% 79.2% 76.0% 71.2% 71.17%
Prioritize diversity w/ order 83.4% 50.8% 81.1% 75.0% 71.6% 72.37%
w/o order 80.5% 46.1% 80.1% 76.0% 65.9% 70.11%
@@
4.2 Main Results and Analysis
We compare the performance of our approach with other ICL methods. All the comparison rusults are tabulated in Table 1. Experimental results show that compared with other ICL methods, we achieve the best performance on SVAMP, AQuA, Gsm8k and StrategyQA. Overall, our method improves the average accuracy of all benchmarks by 2.24%. This result shows that our method effectively improves the model’s reasoning performance. To further demonstrate the effectiveness of the method, we conducted multiple sets of experiments for illustration.
@@ 4.2.1 Analysis of Example Selection and Ordering
The complete experimental results are shown in Table 2, and through analysis, we have made the following observations:
@@ 5.1 In-Context Learning
GPT-3 Brown et al. (2020) exhibited few-shot and zero-shot learning abilities during the pretraining phase. CoT Wei et al. (2022) designed several fixed demonstration examples manually as in-context information, inspired further research on ICL Yao et al. (2024).

### v2

@@ Abstract
In-context learning (ICL) can significantly enhance the complex reasoning capabilities of large language models (LLMs), with the key lying in the selection and ordering of demonstration examples. Previous methods typically relied on simple features to measure the relevance between examples. We argue that these features are not sufficient to reflect the intrinsic connections between examples. In this study, we propose a curriculum ICL strategy guided by problem-solving logic. We select demonstration examples by analyzing the problem-solving logic and order them based on curriculum learning. Specifically, we constructed a problem-solving logic instruction set based on the BREAK dataset and fine-tuned a language model to analyze the problem-solving logic of examples. Subsequently, we selected appropriate demonstration examples based on problem-solving logic and assessed their difficulty according to the number of problem-solving steps. In accordance with the principles of curriculum learning, we ordered the examples from easy to hard to serve as contextual prompts. Experimental results on multiple benchmarks indicate that our method outperforms previous ICL approaches in terms of performance and efficiency, effectively enhancing the complex reasoning capabilities of LLMs. Our project will be released at https://github.com/maxuetao/CurriculumICL
@@ 1 Introduction
Large language models (LLMs) Ouyang et al. (2022); Ye et al. (2023); Bahrini et al. (2023) can rapidly acquire new capabilities through in-context learning (ICL) to solve many new tasks Wies et al. (2024); Xu et al. (2024), and can be extended through chain of thought (CoT) Wei et al. (2022) to solve many tasks that require complex reasoning Zhang et al. (2023); Zhao and Zhang (2024); Hou et al. (2025). Researchers believe that through ICL, LLMs can implicitly learn the problem-solving patterns demonstrated in contextual examples and apply them to new tasks Bhattamishra et al. (2023); Dai et al. (2023). This means that LLMs have the ability to learn and apply problem-solving patterns on the spot from given examples.
In recent years, supervised fine-tuning (SFT) methods Dong et al. (2023); Du et al. (2025a) and reinforcement learning methods Du et al. (2023); Guo et al. (2025); Du et al. (2025b) have been able to significantly enhance the abilities of LLMs through training. Despite this, due to the unique characteristic of ICL that it can enhance problem-solving capabilities without training, it still holds value as significant as the methods mentioned above, especially when facing the need to reduce costs or quickly apply to new tasks. Relevant work Hsieh et al. (2023) has already shown that LLMs possess a wealth of basic knowledge and fundamental capabilities that can be effectively activated through a small number of examples. Particularly, LIMO Ye et al. (2025) fine-tuned a large language model with only a few hundred examples and achieved results that are close to or even on par with the current state-of-the-art reinforcement learning optimization inference. Therefore, we believe that the ICL capabilities of current LLMs are still far from being fully realized. There is a need to design better prompts to effectively enhance the effectiveness of ICL.
@@
In this work, we innovatively propose a problem-solving logic guided curriculum ICL method, which constructs the optimal ICL prompt for the query based on problem-solving logic. The Question Decomposition Meaning Representation (QDMR) Wolfson et al. (2020) decomposes complex problems into several sub-questions for solving and formalizes these sub-questions with 13 custom "operations", which we refer to as problem-solving logic. Figure 1-(a) shows an example of problem decomposition and transformation into problem-solving logic. Although it cannot directly solve the problem, the problem-solving logic describes the steps required for solving and the order of these steps in formal language. Therefore, it can accurately measure the intrinsic connections between examples and construct a sequence of demonstration examples that are conducive to problem-solving. Figure 1-(b) shows an example of curriculum ICL. We select examples with similar problem-solving logic, which can help LLMs learn how to solve similar problems. Subsequently, we measure the difficulty of these examples by the number of problem-solving steps. The greater the number of steps, the more reasoning steps are involved, meaning the problem is more difficult to solve. Relying on the principles of curriculum learning, we order these examples from easy to hard to serve as the final in-context prompt.
@@ 3.2.2 Demonstration Examples Ordering
Llama3-8B
Random Random Random 76.5 46.5 73.8 75.8 65.1 67.53
VoteK KNN Similarity 74.9 44.9 76.7 75.4 69.0 68.19
PromptSO PCA Eigenvalue 77.3 43.7 77.7 75.6 67.7 68.40
AutoCoT K-means Similarity 77.5 47.2 75.3 76.0 71.2 69.44
CoT + Fewshot Fixed Fixed 80.5 44.5 79.4 75.1 69.4 69.79
SA-ICL KNN Entropy 78.8 47.6 77.9 78.5 66.8 69.95
AL-ICL KNN Similarity 80.8 45.7 78.2 77.9 68.1 70.13
Ours PSL Curriculum 83.4 50.8 81.1 75.0 71.6 72.37
Llama3-70B
Random Random Random 85.9 69.7 91.6 81.7 80.3 81.84
VoteK KNN Similarity 86.2 67.3 92.3 81.6 82.5 81.98
PromptSO PCA Eigenvalue 86.8 73.6 90.1 81.0 77.3 81.76
AutoCoT K-means Similarity 89.2 65.0 92.5 81.4 74.2 80.46
CoT + Fewshot Fixed Fixed 88.8 67.7 91.6 81.4 82.5 82.40
SA-ICL KNN Entropy 85.7 67.3 92.3 81.1 82.5 81.78
AL-ICL KNN Similarity 87.2 68.5 91.7 81.2 84.2 82.56
Ours PSL Curriculum 90.5 70.1 92.6 81.8 85.2 84.04
Qwen2.5-7B
Random Random Random 87.3 74.8 87.2 83.4 65.0 79.54
VoteK KNN Similarity 86.2 79.9 84.6 83.5 68.1 80.46
PromptSO PCA Eigenvalue 85.6 76.4 90.5 82.7 70.3 81.11
AutoCoT K-means Similarity 87.1 69.7 90.4 83.6 69.7 80.10
CoT + Fewshot Fixed Fixed 90.2 76.4 89.8 83.8 63.8 80.79
SA-ICL KNN Entropy 87.7 71.7 88.7 83.3 69.0 80.07
AL-ICL KNN Similarity 88.4 77.6 89.5 82.5 67.7 81.13
Ours PSL Curriculum 92.3 78.3 90.0 84.6 71.2 83.28
Table 1: The table presents a comparison of experimental results across different benchmarks using Llama3-8B, Llama3-70B and Qwen2.5-7B, demonstrating the accuracy contrast between various ICL methods. Avg represents the average accuracy across the different benchmarks. The best and second-best performances are highlighted in bold and underlined, respectively.
@@ 4.1 Experimental Setup
We evaluate the effectiveness of our method on the Llama3-8B, Llama3-70B and Qwen2.5-7B Qwen et al. (2025). For each benchmark, we select demonstration examples from its training set to form prompt information to evaluate each test set data. For the SVAMP dataset, we adopted the same evaluation strategy as in previous work Patel et al. (2021), using ASDiv-a Miao et al. (2020) and MAWPS Koncel-Kedziorski et al. (2016) together as the training set. To ensure a fair comparison, the number of selected examples for all baselines is based on the settings in CoT Wei et al. (2022) for different benchmarks, and the number of examples selected by our method does not exceed that of the baselines.
4.2 Main Results and Analysis
We compare the performance of our approach with other ICL methods. All the comparison results are tabulated in Table 1. For Llama3-8B, compared with other methods, we achieve the best performance on SVAMP, AQuA, Gsm8k and StrategyQA. The average accuracy across all benchmarks is improved by 2.24%. For Llama3-70B, we achieve the best performance on SVAMP, Gsm8k, ComSenQA and StrategyQA. The average accuracy across all benchmarks is improved by 1.48%. For Qwen2.5-7B, we achieve the best performance on SVAMP, ComSenQA and StrategyQA. The average accuracy across all benchmarks is improved by 2.15%. Overall, our proposed method demonstrates excellent performance across different parameter scales and different models, which shows the effectiveness and generalizability of the method.
To further demonstrate the effectiveness of the method, we have conducted validation and analysis of the key strategies of the method.
@@ 4.2 Main Results and Analysis
Prioritize simplicity w/ order 82.3 47.6 79.5 75.5 69.0 70.79
w/o order 82.5 47.2 78.8 76.1 68.1 70.55
Prioritize difficulty w/ order 81.8 44.9 77.9 76.6 67.7 69.77
w/o order 81.6 46.1 79.6 77.0 67.2 70.29
Select Randomly w/ order 81.3 50.6 80.2 76.1 70.3 71.70
w/o order 80.9 48.6 79.2 76.0 71.2 71.17
Prioritize diversity w/ order 83.4 50.8 81.1 75.0 71.6 72.37
w/o order 80.5 46.1 80.1 76.0 65.9 70.11
@@ 4.2.1 Analysis of Example Selection and Ordering
We conducted experiments on Llama3-8B, and the results are shown in Table 2. Through analysis, we have made the following observations:
@@ 5.1 In-Context Learning
GPT-3 Brown et al. (2020) exhibited few-shot and zero-shot learning abilities during the pretraining phase. CoT Wei et al. (2022) designed several fixed demonstration examples manually as in-context information, inspired further research on ICL Li et al. (2025a); Yao et al. (2024); Li et al. (2025b).
@@ Acknowledgements
This project was supported by the Natural Science Foundation of China (Project Number 62437001).